OpenStudy (anonymous):

the length of a rectangle is twice the width the perimeter is 48 in find the demensions of the rectangle

OpenStudy (ajprincess):

It is given that the length of the rectangle is twice of the width. so if the length is l and width is w then l=2w. the perimeter of a rectangle is found by the sum of lengths of all sides. ie perimeter=l+w+l+w 48=2l+2w but l=2w so 48=2(2w)+2w 48=4w+2w Find w. then plug in that value in l=2w to find l. Does that help?